# Don’t Pay Attention

### v1

@@ Appendix I Learning Rate: To Decay or Not to Decay?
Appendix M Additional Long-Range Benchmark Results
@@
In Section 3.4, we presented results for Avey, Transformer++, Mamba, and RWKV-7 under the medium parameter regime ($\sim$500M parameters) on the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38], which is designed to evaluate models’ abilities to handle long-range dependencies. The S-NIAH benchmark, along with two of its common variants—S-NIAH-1 and S-NIAH-2—was described in detail in Section 3.4. In this section, we extend our analysis by reporting results under two additional model regimes, small ($\sim$150M parameters) in Fig. 7 and large ($\sim$1.5B parameters) in Fig. 8. Akin to the experiment in Section 3.4, Transformer++, Mamba, and RWKV-7 were trained with a context window of 2048 tokens, while Avey was trained with a shorter window of only 512 tokens.
@@
In this paper, we kept Avey’s context window fixed at 512 tokens across all parameter regimes. All tuning experiments related to sequence length, split size, and top $k$ splits (see Section G) were conducted exclusively using the small model size. It is plausible that with larger capacity, Avey could more effectively leverage longer sequences by retrieving and contextualizing a larger set of relevant tokens while filtering out less informative ones, thereby enhancing contextual representations and further boosting performance. Investigating the relationship between sequence length and model size in Avey is an interesting direction for future work.
Appendix N Complexity Analysis
@@
During inference, the time complexities of the enricher, contextualizer, and fuser remain unchanged. However, the ranker’s analysis slightly changes, as at each time step $t$ (i.e., upon predicting a new token), the current split is compared against all previous splits. More precisely, at each time step $t$, the current split—denoted as split $i$ and incrementally filled as tokens are generated—is compared against all $i-1$ preceding splits, each consisting of $S$ tokens. Consequently, the cost of comparing $t$ tokens in split $i$ (with $t\leq S$) against $S$ tokens in a previous split is $\mathcal{O}(t\cdot S\cdot d)$.
@@
Appendix O Related Work
Recurrent Neural Networks (RNNs) [19, 79] are designed to process sequential data by capturing temporal dependencies, making them well-suited for tasks where input order is essential. However, their cyclical nature limits their potential for parallel computation and exposes them to vanishing and exploding gradient problems. As a result, they typically struggle to effectively learn long-range dependencies. While architectures like Long Short-Term Memory (LSTM) [34] and Gated Recurrent Units (GRU) [8] mitigate these gradient-related issues, they remain slow to optimize and challenging to scale due to retaining RNN’s core recurrent and sequential structure.
In contrast, the Transformer [98] employs a self-attention mechanism to process each sequence of tokens simultaneously. More precisely, it promotes two key design principles: (1) a recurrent-free architecture, which enables parallel computation of token embeddings, while still capturing their order through positional encodings, and (2) a multi-head self-attention approach, which facilitates cross-token interactions to further enrich the expressiveness of embeddings. These innovations make the Transformer highly effective, as well as parallelizable and efficient to train. However, they also cause its computational and memory requirements to scale quadratically with sequence length, making it expensive and less efficient for very long sequences.
To address the Transformer’s quadratic computation and memory costs, a wide array of approaches have been proposed, including linear attention [50, 48, 9, 71, 106], sparse or local attention [104, 7, 67], context compression [74, 99, 90, 78], and modified attention computations [93, 100, 92], to mention just a few. Notably, the Attention-Free Transformer (AFT) [106] offers a linear drop-in replacement for the quadratic self-attention mechanism. In particular, it weights key and value vectors using learned positional biases and integrates them with query vectors via element-wise multiplication. As such, it eliminates the need to compute and store the costly attention matrix while still preserving global query-value interactions—without requiring architectural modifications or additional tuning. Furthermore, AFT introduces variants such as AFT-local and AFT-conv, which leverage local attention patterns to reduce parameter count and further improve computational and memory efficiency.
RWKV-4 [68] (the first 3 versions were experimental [52]) capitalizes on AFT and suggests combining the strengths of both Transformers and RNNs. To elaborate, unlike Transformers and akin to RNNs, it does not process each input token solely based on its own embedding, but rather as a weighted sum of its embedding and that of the preceding one. To the contrary of traditional RNNs and similar to Transformers, it adopts self-attention, but an extended version of it, namely, that of AFT. This hybrid approach allows RWKV-4 to maintain some of the modeling capabilities of RNNs, while benefiting from the parallelization and scalability features of Transformers.
More precisely, RWKV-4 extends AFT in two distinct ways: (1) it introduces an additional parameter to handle each current token independently, and (2) it implements a per-time-step decay mechanism that selectively removes older content from the recurrent hidden state in a data-dependent manner. This decay mechanism addresses a central limitation of linear attention, which pertains to the lack of a systematic way to discard outdated information [83, 102].
RWKV-5 [69] enhances RWKV-4’s architecture and learning decay mechanism by replacing traditional vector-valued states with more expressive multi-head, matrix-valued ones. Moreover, it reconfigures receptive states, incorporates supplementary gating mechanisms, and dynamically learns the linear interpolation between the current and preceding token embeddings instead of relying on pre-defined hyperparameters.
RWKV-6 [69] promotes a new application of low-rank adaptation functions [39, 52]. Specifically, it makes the linear interpolation between the current and preceding tokens data-dependent to improve the selectivity of the model in retaining and discarding information. Additionally, it replaces the static decay mechanism with a dynamic one, allowing each element in the decay vector to fluctuate independently over time in response to the input.
The decay strategies of RWKV-4, RWKV-5, and RWKV-6 still cannot remove values stored at specific keys. DeltaNet [83] overcomes this drawback by partially replacing the values stored at current keys with equivalent new values, enabling models to erase outdated memories and include up-to-date ones on a per-key basis. However, it only allows a fixed scalar fraction of a value to be replaced from a state via an in-context learning rate parameter, thus demonstrating rigidity in adapting to varying data contexts.
RWKV-7 [70] builds upon the principles of DeltaNet and introduces a vector-valued in-context learning rate instead of a scalar-valued one. This allows selective replacement of state data on a channel-wise basis. Furthermore, RWKV-7 employs a vector-valued decay mechanism and uses additional low-rank projections to optimize the trade-off between the number of parameters, computational efficiency, and downstream performance. Lastly, it incorporates Value Residual Learning [110], which improves the propagation of initial local information via utilizing a residual connection between the value vectors of the current layer and those of the first layer prior to the attention operation, resulting in enhanced language modeling performance.
Most recently, RWKV-X [37] proposed combining the strengths of RWKV and sparse attention, drawing inspiration from Mixture of Block Attention (MoBA) [57]. In particular, RWKV-X restricts each query to attend only to a small, relevant subset of the input, thus reducing computational cost and facilitating the modeling of longer-range dependencies. More precisely, rather than allowing each token to attend to every other token in the sequence (as in traditional self-attention), it constrains each token’s attention to a limited subset (hence, making it sparse), while maintaining the coupling between the input sequence and context window.
Similar to RWKV, RetNet [91] adopts linear attention and promotes a hybrid approach that blends Transformer- and RNN-like representations, yet with a decay-based memory unit. Specifically, it divides the input sequence into chunks, wherein the Transformer-like parallel representation is applied. Additionally, it enables propagating information sequentially across chunks using the RNN-like representation. Lastly, it uses multiple attention heads, each governed by a distinct decay rate, and replaces LayerNorm [1] with GroupNorm [101].
Although linear attention has been proposed as a promising alternative to quadratic softmax attention [48, 9, 47, 71], existing implementations of it are in practice slower than optimized versions of softmax attention [14, 12, 102]. From an accuracy standpoint, linear attention generally underperforms conventional softmax attention, sometimes by a significant margin in language modeling [47, 102].
To this end, and in light of the exponentially growing complexity associated with overcoming the limitations of Transformer-based architectures, there has been a renewed interest in RNN-based alternatives in recent years. Notably, Structured State Space Sequence (S4) models [27, 28], inspired by the classical state space models (SSMs) [44], have emerged as a promising paradigm for sequence modeling. These models describe the temporal evolution of a system using differential equations, offering a continuous-time formulation of dynamics, and can be viewed as generalized versions of RNNs.
An SSM as a concept has a broad meaning, which simply refers to the notion of any recurrent process with a latent state [25]. From this perspective, the RNN-like linear attention model proposed and formulated by [48] can be interpreted as a degenerate linear SSM. Interestingly, this justifies the usage of a decay factor in RetNet and RWKV, especially that a decay term (or a forget gate) has been shown to be crucial in RNNs [34, 97, 8].
Numerous variants of SSMs [27, 29, 30, 51, 58, 65, 87] have demonstrated strong performance across a range of domains, including audio and vision [24, 63, 81]. Nonetheless, these variants have struggled with language modeling, often lagging behind Transformers by several points in perplexity [27].
From an efficiency standpoint, however, SSMs have shown encouraging results in language modeling. For instance, S4 [27, 28], a prominent SSM, converts the continuous-time state update equation of SSMs into a discrete form, hence, enabling parallel sequence modeling. Moreover, it utilizes the HiPPO (High-Order Polynomial Projection Operator) initialization [26], which alleviates the vanishing gradient problem and facilitates processing longer sequences.
Another example of SSMs is H3 [20], which improves language modeling by allowing both, the recall of earlier tokens and token-wise comparisons within a sequence. H3 extends S4 by suggesting a state-passing algorithm that enhances computational efficiency on modern hardware. This advancement reduces the hardware-related barriers that have traditionally limited the scalability of SSM-based architectures.
Hyena [72] capitalizes on H3 by replacing its S4 layer with an MLP-parameterized global convolution [77]. S5 [87] proposes using parallel scan [59] to parallelize S4. Liquid S4 [31] augments S4 with an input-dependent state transition matrix, computed convolutionally in the frequency domain (which is computationally efficient) and mapped back to the time domain using an inverse Fourier transformation. SGConv [51], LongConv [21], MultiresConv [86], and Toeplitz Neural Network [73] all focus on the convolutional representation of S4 as well, aiming to enhance its efficiency [25].
Most recently, Mamba [25] introduced a new class of SSMs known as selective SSMs, specifically designed to improve the performance of language modeling tasks. Mamba addresses a key limitation in SSMs, namely, their inability to selectively process inputs in an input-dependent manner (i.e., focus on or ignore specific parts of the input sequence). Consequently, it makes the SSM parameters input-dependent, but introduces a technical challenge since traditional SSMs are inherently designed to be time- and input-invariant to ensure computational efficiency. To overcome this challenge, Mamba proposes a hardware-efficient parallel scan (or prefix sum) algorithm [6], which enables recurrent-style computation without explicitly materializing the expanded state. This design precludes costly I/O operations across GPU memory hierarchies and accelerates both, training and inference.
Tri Dao and Albert Gu [13] argue that various approaches to operating SSMs can be reframed as matrix multiplication algorithms involving a specific class of structured matrices known as semiseparable matrices. They further leverage the language of tensor contractions to prove the recurrent formulation of linear attention as proposed by [48], before generalizing it to a new family of structured masked attention (SMA).
Subsequently, Tri Dao and Albert Gu demonstrated that SSMs and Transformers are fundamentally connected, governed by the mathematical framework of semiseparable matrices and SMA. Additionally, they developed a rich state space duality (SSD) framework of theoretical connections between SSMs and various forms of attention. This framework facilitated the design of Mamba-2, an extended version of Mamba, which aims to improve its efficiency (not performance). Mamba-2 utilizes a scalar data-dependent gating mechanism (like the ones proposed by [71, 91, 3]), which enables transforming its recurrent structure into a matrix-multiply form, thus allowing for efficient execution on tensor cores and better support for larger hidden state sizes.
The strategy of the SSD framework mirrors that of linear attention [48], which established a connection between autoregressive attention mechanisms and linear RNNs via showing an equivalence between "dual forms" of quadratic kernelized attention and a specific type of linear recurrence. Conceptually, the SSD framework seeks to transfer algorithmic and systems-level optimizations originally developed for Transformers to the realm of SSMs. Its overarching goal is to enable the development of architectures that outperform Transformers, while scaling more efficiently with sequence length.
Finally, several works, including Tolstikhin et al. [95], Melas-Kyriazi [60], Touvron et al. [96], and Ding et al. [18], among others, have questioned the necessity of self-attention, particularly in the context of Vision Transformers. In contrast, Liu et al. [53] introduced gMLP, an MLP-based alternative to BERT-style Transformers [16] that (partially) eliminates self-attention but ultimately underperforms average performance on downstream NLP tasks. gMLP encompasses channel (hidden) and spatial (cross-token) projections with multiplicative gating and static parameterization. Its gating mechanism is reminiscent of Gated Linear Units (GLUs) [15, 84, 100], as well as earlier architectures such as Highway Networks [88] and LSTM-RNNs [34]. A key distinction, however, is that gMLP applies gating on the spatially projected dimension and not the hidden one. The gated embedding-wise neural network in Avey’s contextualizer draws inspiration from gMLP.
Unlike all previously mentioned models, Avey entirely abandons self-attention and recurrence, introducing a new architecture composed of a ranker and a dynamically parameterized neural processor. The ranker identifies the most relevant tokens for contextualization, while the neural processor contextualizes them data-dependently. This design decouples sequence length from context width, enabling efficient processing of arbitrarily long sequences without diminishing the influence of distant yet important tokens.
At the core of Avey’s architecture is a weighted-selective-split interaction mechanism, which filters out irrelevant tokens beyond the context window and enables direct interactions only with relevant ones, thus preserving their influence irrespective of sequence length. In addition, Avey employs a partial-embedding bypassing technique that retains a portion of each token’s raw, distinctive features before fusing them with its contextualized ones through a neural network. This technique boosts the performance of Avey (as shown in Appendix K) and might help mitigate issues such as entropy collapse [107] and over-smoothing [109, 85], especially at large-scale, when the depth of the model is increased significantly.
Appendix P Limitations
@@
The scope of our work is limited to textual data and does not involve other modalities such as images, audio, or genomics. Additionally, our evaluation of Avey is restricted to standard autoregressive language modeling, benchmarking it against popular open-source architectures using both pretraining metrics (perplexity) and zero-shot evaluations on established NLP benchmarks. As a result, we do not investigate Avey’s ability to construct bidirectional contextualized word representations, as done in BERT [16]. We leave this for future work. Finally, the paper focuses solely on effectiveness rather than efficiency. While we provide a complexity analysis showing that Avey exhibits quadratic training time like Transformers, our current implementation is slower. As such, further engineering efforts are required to optimize it.

### v2

@@ Appendix B Additional Short-Range Benchmark Results
Appendix C Additional Long-Range Benchmark Results
@@
In Section 3.5, we presented results for Avey, Transformer++, Mamba, and RWKV-7 under the medium parameter regime ($\sim$500M parameters) on the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [hsieh2024ruler], which is designed to evaluate models’ abilities to handle long-range dependencies. The S-NIAH benchmark, along with two of its common variants—S-NIAH-1 and S-NIAH-2—was described in detail in Section 3.5. In this section, we extend our analysis by reporting results under two additional model regimes, small ($\sim$150M parameters) in Fig. 7 and large ($\sim$1.5B parameters) in Fig. 8. Akin to the experiment in Section 3.5, Transformer++, Mamba, and RWKV-7 were trained with a context window of 2048 tokens, while Avey was trained with a shorter window of only 512 tokens.
@@
In this paper, we kept Avey’s context window fixed at 512 tokens across all parameter regimes. All tuning experiments related to sequence length, split size, and top $k$ splits (see Section 3.2.6) were conducted exclusively using the small model size. It is plausible that with larger capacity, Avey could more effectively leverage longer sequences by retrieving and contextualizing a larger set of relevant tokens while filtering out less informative ones, thereby enhancing contextual representations and further boosting performance. Investigating the relationship between sequence length and model size in Avey is an interesting direction for future work.
Appendix D Complexity Analysis
@@
During inference, the time complexities of the enricher, contextualizer, and fuser remain unchanged. However, the ranker’s analysis slightly changes, as at each time step $t$ (i.e., upon predicting a new token), the current split is compared against all previous splits. More precisely, at each time step $t$, the current split—denoted as split $i$ and incrementally filled as tokens are generated—is compared against all i1i-1 preceding splits, each consisting of $S$ tokens. Consequently, the cost of comparing $t$ tokens in split $i$ (with $t\leq S$) against $S$ tokens in a previous split is $\mathcal{O}(t\cdot S\cdot d)$.
@@
Appendix E Is the Ranker a RAG Component?
The ranker is an internal component of Avey that operates within the input sequence, selecting among its splits for more effective contextualization. It does not query external corpora or indexes, introduces no retrieval I/O or freshness dependencies, and adds no retrieval latency. Its role is architectural, that is, to allocate Avey’s internal contextual budget and decouple context width from sequence length so that Avey can fully contextualize sequences far beyond its training window.
By contrast, Retrieval-Augmented Generation (RAG) [lewis2020rag] augments a model with external (non-parametric) knowledge via a retriever, ranging from classic BM25 [robertson2009bm25] and dense passage retrieval (DPR) [karpukhin2020dpr] to system-level designs such as REALM [guu2020realm], RETRO [borgeaud2022retro], and MacRAG [lim2025macrag], among others. RAG aims to: (1) improve factuality by grounding outputs in retrieved documents, (2) make models updatable by reflecting new information without retraining, and (3) reduce compute for long-context tasks by moving knowledge out of weights.
As such, the two mechanisms are orthogonal. The ranker allocates the model’s internal contextual budget over the given sequence, whereas RAG changes the evidence set by importing out-of-sequence content. They can be composed (RAG can be layered atop Avey, as it is with the Transformer) but one does not subsume the other.
Appendix F Neural Contextualization vs. Attention
Avey’s contextualizer is an embedding-wise neural network that dispenses with attention. In Appendix 3.3, we replaced it with standard self-attention and observed a $4.6\%$ increase in perplexity alongside a $2.1\%$ decline in average task performance, underscoring its central role in Avey’s architecture.
Formally, the contextualizer is defined in Equation 2, repeated below for convenience:
$\mathbf{c}(\mathbf{Z}_{t})=\mathbf{Z}_{tl}\ \odot\ \sigma\!\Big(\big(\mathbf{V}\ \odot\ \mathcal{N}(\mathbf{Z}_{tr})\,\mathcal{N}(\mathbf{Z}_{tr})^{\top}\big)\,\mathbf{Z}_{tr}\ +\ \mathbf{b}^{\prime}\Big).$
Let $\mathbf{S}:=\mathcal{N}(\mathbf{Z}_{tr})\,\mathcal{N}(\mathbf{Z}_{tr})^{\top}$. The product $(\mathbf{V}\!~\odot\!~\mathbf{S})\mathbf{Z}_{tr}$ yields a content-dependent signal that is passed through a pointwise nonlinearity and used to gate $\mathbf{Z}_{tl}$ elementwise, producing a bounded, feature-wise modulation rather than a mixture over values. By contrast, self-attention computes a row-stochastic convex combination of value vectors after a Q/K split and softmax normalization. Evidently, our formulation departs away from both the softmax and the Q/K/V decomposition, whereby weights are neither constrained to be nonnegative nor to sum to one, and the output acts as a gate on carrier features (i.e., $\mathbf{Z}_{tl}$) rather than a convex average of value vectors.
The contextualizer also differs fundamentally from linear attention [choromanski2021performer, katharopoulos2020transformers, wang2020linformer, beltagy2020longformer, sun2023retentive]. Linear-attention variants obtain near-linear complexity by exploiting an associative kernel factorization that permits reordering and prefix accumulation, typically of the form $\phi(\mathbf{Q})\big(\phi(\mathbf{K})^{\top}\mathbf{V}_{\!\text{val}}\big)$. Equation 2 does not admit such reordering. In fact, the Hadamard coupling ($\mathbf{V}\!~\odot\!~\mathbf{S}$) breaks the algebraic associativity required to push multiplications across terms, and the normalization $\mathcal{N}(\cdot)$ is neither linear nor guaranteed nonnegative, precluding the kernel tricks used to approximate softmax attention with associative feature map functions (e.g., ReLU and Exp). Lastly, we note that the contextualizer remains quadratic (not linear) in sequence length.
For similar reasons, Equation 2 cannot be reformulated as a finite-state RNN under an autoregressive mask. Let $\mathbf{S}_{t}=\mathcal{N}(\mathbf{Z}_{tr}^{\leq t})\,\mathcal{N}(\mathbf{Z}_{tr}^{\leq t})^{\top}$. The update at step t+1t\!+\!1 depends on the full pairwise matrix $(\mathbf{V}\!~\odot\!~\mathbf{S}_{t})$, that is, on all position-specific interactions among the past tokens after data-dependent normalization. Because the learned weight matrix $\mathbf{V}$ introduces position-dependent multiplicative couplings, there is no time-invariant transition $h_{t+1}=f(h_{t},x_{t+1})$ with a fixed-dimensional sufficient statistic $h_{t}$ that exactly summarizes $(\mathbf{V}\!~\odot\!~\mathbf{S}_{t})$. In particular, the required weights vary across positions and must be recomputed, so any streaming recurrence would either approximate by tying/averaging $\mathbf{V}$ or maintain $\mathcal{O}(t)$ state. Therefore, an exact finite-state RNN equivalence is unavailable.
Empirically, $\mathbf{V}$ performs most of the heavy-lifting in Avey, while $\mathbf{S}$ primarily induces selectivity, dynamically emphasizing or suppressing interactions conditioned on the input, echoing the selectivity principle advocated in recent sequence models [gu2023mamba]. An ablation in Appendix 3.3 shows that including $\mathbf{S}$ delivers a consistent, albeit modest, gain by making the neural processor’s parametrization input-adaptive.
Putting everything together, these distinctions (i.e., gating rather than mixing, non-associative pairwise modulation rather than kernel-factorizable operations, and explicit quadratic interactions), explain both the theoretical departure from self-attention and linear attention and the observed empirical contribution of the contextualizer within Avey.
Appendix G Design Rationale
We designed Avey around clear functional roles for its core modules. Below, we outline some of the guiding intuitions and how they inform its architecture.
Enricher:
A substantial body of evidence indicates that much of a language model’s knowledge is stored in feed-forward sublayers and accessed through non-linear feature interactions (e.g., [geva2021transformer_ffn_memory]). The enricher is designed accordingly. It serves both as the primary repository of parametric knowledge and as a mechanism for intra-embedding interactions, enabling higher-order, non-linear composition of features within each embedding. This improves expressivity by allowing features to modulate and refine one another in a context-aware manner.
Contextualizer:
The contextualizer operates as an embedding-wise neural network such that each neuron forms a weighted sum over input embeddings with learned coefficients (see Equation 2). To introduce input-dependent selectivity (as in [gu2023mamba]), we augment these static weights with a cosine-similarity term that produces a second, data-driven set of weights (the two are combined multiplicatively via a Hadamard product). This dynamic modulation improves behaviors such as copying and induction by strengthening interactions that are semantically relevant to the current input. The split-and-gate structure follows established gated designs in gMLP [liu2021pay] and GLU variants [dauphin2017language, shazeer2020glu, wu2019pay].
Partial Embedding Bypassing:
The enricher’s output is partitioned into two streams, one is passed to the contextualizer and the other is bypassed and fed directly to the fuser. The bypassed part plays two complementary roles. First, it provides a strong residual path that preserves signal and stabilizes optimization by improving gradient flow within each Avey layer. Second, it supplies additional non-linear capacity in the downstream feed-forward fuser, complementing the contextualizer’s primarily linear mixing across embeddings. This balance between context-aware and context-invariant processing yields richer, more diverse representations.
Fuser:
The fuser (a position-wise feed-forward network) learns how to combine the contextualized and bypassed streams and then projects the result back to the model’s embedding dimension, ensuring compatibility with residual pathways across layers. As a feed-forward network (FFN), it also contributes to storing and accessing parametric knowledge learned during training, analogous to FFN roles in Transformers.
Appendix H Limitations
@@
The scope of this work is limited to textual data and does not extend to other modalities such as images or audio, nor to specialized domains like genomics. Our evaluation of Avey focuses on standard autoregressive language modeling, benchmarking against popular open-source architectures using both pretraining metrics (perplexity) and zero-shot performance on established NLP benchmarks. Consequently, we do not investigate Avey’s ability to produce bidirectional contextualized representations as in BERT [devlin2019bert]. We leave this direction to future work. Finally, the paper emphasizes effectiveness rather than efficiency. While our complexity analysis indicates quadratic training time similar to Transformers, our current implementation is slower in practice, and additional engineering is needed to optimize it.